# Detection and Recovery of Hidden Submatrices

### v1

@@ 4.5.3 Correlated recovery using the peeling estimator
where we define ${\cal S}_{1}(\mathsf{K})\triangleq\sum_{(i,j)\in\mathsf{K}}\mathsf{X}_{ij}^{(1)}$, for $\mathsf{K}\in{\cal K}_{k,1,n}^{\mathsf{con}}$. Under the planted model, ${\cal S}_{1}(\mathsf{K})\sim{\cal N}(\lambda\langle{\mathsf{K},\mathsf{K}^{\star}_{1}}\rangle,k^{2})$. Hence, the distribution of ${\cal S}_{1}(\mathsf{K})$ depends on the size of the overlap of $\mathsf{K}$ with $\mathsf{K}_{1}^{\star}$. To prove that reconstruction is possible, we compute in the planted model the probability that ${\cal S}_{1}(\mathsf{K})>{\cal S}_{1}(\mathsf{K}_{1}^{\star})$, given that $\mathsf{K}$ has overlap $\langle{\mathsf{K},\mathsf{K}_{1}^{\star}}\rangle=\omega$ with the planted partition, and argue that this probability tends to zero whenever the overlap is small enough. Taking the union bound over every $\mathsf{K}$ with overlap at most $\omega$ gives
$\mathbb{P}\left(\max_{\langle{\mathsf{K},\mathsf{K}_{1}^{\star}}\rangle\leq\omega}{\cal S}_{1}(\mathsf{K})>\lambda k^{2}-\sqrt{2k^{2}\log n}\right)\leq$ (119)
$\hskip 22.76228pt=n\cdot\max_{\langle{\mathsf{K},\mathsf{K}_{1}^{\star}}\rangle\leq\omega}\exp\left(-\frac{\left[\lambda(k^{2}-\langle{\mathsf{K},\mathsf{K}_{1}^{\star}}\rangle)-\sqrt{2k^{2}\log n}\right]^{2}}{2k^{2}}\right)$ (120)
$\hskip 22.76228pt=n\cdot\max_{\langle{\mathsf{K},\mathsf{K}_{1}^{\star}}\rangle\leq\omega}\exp\left(-\left[\frac{\lambda}{\sqrt{2k^{2}}}(k^{2}-\langle{\mathsf{K},\mathsf{K}_{1}^{\star}}\rangle)-\sqrt{\log n}\right]^{2}\right)$ (121)
$\hskip 22.76228pt\leq\exp\left(\log n-\left[\frac{\lambda}{\sqrt{2k^{2}}}(k^{2}-\omega)-\sqrt{\log n}\right]^{2}\right).$ (122)
By the assumption that $\lambda>\frac{2\sqrt{2\log n}}{k}$, it follows that there exists a fixed constant $\epsilon>0$ such that $(1-\epsilon)\lambda>\frac{2\sqrt{2\log n}}{k}$. Hence, setting $\omega=k^{2}\epsilon$ in the last displayed equation, we get
$\mathbb{P}\left(\max_{\langle{\mathsf{K},\mathsf{K}_{1}^{\star}}\rangle\leq\omega}{\cal S}_{1}(\mathsf{K})>\lambda k^{2}-\sqrt{2k^{2}\log n}\right)$
$\hskip 22.76228pt\leq\exp\left(\log n-\left[\frac{\lambda\sqrt{k^{2}}}{\sqrt{2}}(1-\epsilon)-\sqrt{\log n}\right]^{2}\right)=e^{-\Omega(n)},$ (123)
@@
$\max_{\langle{\mathsf{K},\mathsf{K}_{1}^{\star}}\rangle\leq k^{2}\epsilon}{\cal S}_{1}(\mathsf{K})<\lambda k^{2}-\sqrt{2k^{2}\log n}.$ (124)
Consequently, we get that the maximum likelihood estimator $\hat{\mathsf{K}}_{1}(\mathsf{X}^{(1)})$ in (20) satisfies $\langle{\hat{\mathsf{K}}_{1},\mathsf{K}_{1}^{\star}}\rangle\geq k^{2}\epsilon$ with high probability. Then, in the second step of the peeling algorithm, we first compute $\mathsf{X}_{ij}^{(2)}$, by setting $[\mathsf{X}_{ij}^{(2)}]_{ij}=-\infty$, for any $(i,j)\in\hat{\mathsf{K}}_{1}$, and $[\mathsf{X}_{ij}^{(2)}]_{ij}=0$, otherwise. Thus, it is clear that in the second step, $\hat{\mathsf{K}}_{2}$ cannot be attained by any set that is $k$-closed to $\hat{\mathsf{K}}_{1}$; indeed, ${\cal S}_{2}(\mathsf{K})=-\infty$, for any set $\mathsf{K}$ that is $k$-closed to $\hat{\mathsf{K}}_{1}$. Therefore, for the relevant sets in the maximization in $\hat{\mathsf{K}}_{2}$, we again have ${\cal S}_{2}(\mathsf{K})\sim{\cal N}(\lambda\langle{\mathsf{K},\mathsf{K}_{2}^{\star}}\rangle,k^{2})$. Accordingly, repeating the exact same arguments above, we obtain that $\langle{\hat{\mathsf{K}}_{2},\mathsf{K}_{2}^{\star}}\rangle\geq k^{2}\epsilon$ with high probability. In the same way, we get that $\langle{\hat{\mathsf{K}}_{\ell},\mathsf{K}_{\ell}^{\star}}\rangle\geq k^{2}\epsilon$ with high probability, for any $\ell\in[m]$. The union bound, then imply that
$\mathbb{P}\left[\bigcup_{\ell=1}^{m}\langle{\hat{\mathsf{K}}_{\ell},\mathsf{K}_{\ell}^{\star}}\rangle<k^{2}\epsilon\right]\leq m\cdot\mathbb{P}\left[\langle{\hat{\mathsf{K}}_{\ell},\mathsf{K}_{\ell}^{\star}}\rangle<k^{2}\epsilon\right]\leq me^{-\Omega(n)}=o(1).$ (125)
@@ 4.6.1 Exact recovery
    We let $\mathsf{K}_{(j-1)\alpha+i}$, for $j=1,2,\ldots,m$ and $i=1,2,\ldots,\alpha$, to be defined the same as $\mathsf{K}_{0}$ but with $\mathsf{S}_{j-1}^{0}$ shifted $i$ columns to the right.
@@
$\leq 2k\frac{\alpha(1+\alpha)}{2}\left[d_{\mathsf{KL}}\left(\mathcal{N}(\lambda,1)||\mathcal{N}(0,1)\right)+d_{\mathsf{KL}}\left(\mathcal{N}(0,1)||\mathcal{N}(\lambda,1)\right)\right]$ (132)
$=\alpha(1+\alpha)k\lambda^{2}.$ (133)
@@
$\lambda^{2}<\frac{\frac{1}{2}\log(1+\mathsf{M})-1}{\alpha(1+\alpha)k}=\frac{\frac{1}{2}\log\mathsf{(1+\alpha m)}-1}{\alpha(1+\alpha)k}.$ (134)
Taking any fixed $\alpha>4$, we get that the minimax error probability is at least half, if $\lambda^{2}<C/k$, for some constant $C>0$, which concludes the proof.

### v2

@@ 4.5.3 Correlated recovery using the peeling estimator
where we define ${\cal S}_{\ell}(\mathsf{K})\triangleq\sum_{(i,j)\in\mathsf{K}}\mathsf{X}_{ij}^{(\ell)}$, for $\mathsf{K}\in{\cal K}_{k,1,n}^{\mathsf{con}}$ and $\ell\in[m]$. Under the planted model, ${\cal S}_{1}(\mathsf{K})\sim{\cal N}(\lambda\langle{\mathsf{K},\mathsf{K}^{\star}}\rangle,k^{2})$. Hence, the distribution of ${\cal S}_{1}(\mathsf{K})$ depends on the size of the overlap of $\mathsf{K}$ with $\mathsf{K}^{\star}$. To prove that reconstruction is possible, we compute in the planted model the probability that ${\cal S}_{1}(\mathsf{K})>\max_{\ell\in[m]}{\cal S}_{1}(\mathsf{K}_{\ell}^{\star})$, given that $\mathsf{K}$ has overlap $\langle{\mathsf{K},\mathsf{K}^{\star}}\rangle=\omega$ with the planted partition, and argue that this probability tends to zero whenever the overlap is small enough. For each $\ell\in[m]$, note that ${\cal S}_{1}(\mathsf{K}_{\ell}^{\star})\sim{\cal N}(\lambda k^{2},k^{2})$, and thus Hoeffding’s inequality implies that ${\cal S}_{1}(\mathsf{K}_{\ell}^{\star})>\lambda k^{2}-\sqrt{2k^{2}\log n}$, with probability at least $1-O(n^{-1})$. Taking the union bound over every $\mathsf{K}$ with overlap at most $\omega$ gives
$\mathbb{P}\left(\max_{\langle{\mathsf{K},\mathsf{K}^{\star}}\rangle\leq\omega}{\cal S}_{1}(\mathsf{K})>\lambda k^{2}-\sqrt{2k^{2}\log n}\right)\leq$ (138)
$\hskip 22.76228pt=n^{2}\cdot\max_{\langle{\mathsf{K},\mathsf{K}^{\star}}\rangle\leq\omega}\exp\left(-\frac{\left[\lambda(k^{2}-\langle{\mathsf{K},\mathsf{K}^{\star}}\rangle)-\sqrt{2k^{2}\log n}\right]^{2}}{2k^{2}}\right)$ (139)
$\hskip 22.76228pt=n^{2}\cdot\max_{\langle{\mathsf{K},\mathsf{K}^{\star}}\rangle\leq\omega}\exp\left(-\left[\frac{\lambda}{\sqrt{2k^{2}}}(k^{2}-\langle{\mathsf{K},\mathsf{K}^{\star}}\rangle)-\sqrt{\log n}\right]^{2}\right)$ (140)
$\hskip 22.76228pt\leq\exp\left(2\log n-\left[\frac{\lambda}{\sqrt{2k^{2}}}(k^{2}-\omega)-\sqrt{\log n}\right]^{2}\right).$ (141)
By the assumption that $\lambda>\frac{\mathsf{C}\sqrt{\log n}}{k}$, with $\mathsf{C}>2+\sqrt{2}$, it follows that there exists a fixed constant $\epsilon>0$ such that $(1-\epsilon)\lambda>\frac{\mathsf{C}\sqrt{\log n}}{k}$. Hence, setting $\omega=k^{2}\epsilon$ in the last displayed equation, we get
$\mathbb{P}\left(\max_{\langle{\mathsf{K},\mathsf{K}^{\star}}\rangle\leq\omega}{\cal S}_{1}(\mathsf{K})>\lambda k^{2}-\sqrt{2k^{2}\log n}\right)$
$\hskip 22.76228pt\leq\exp\left(2\log n-\left[\frac{\lambda\sqrt{k^{2}}}{\sqrt{2}}(1-\epsilon)-\sqrt{\log n}\right]^{2}\right)=e^{-\Omega(n)},$ (142)
@@
$\max_{\langle{\mathsf{K},\mathsf{K}^{\star}}\rangle\leq k^{2}\epsilon}{\cal S}_{1}(\mathsf{K})<\lambda k^{2}-\sqrt{2k^{2}\log n}.$ (143)
Consequently, we get that the maximum likelihood estimator $\hat{\mathsf{K}}_{1}(\mathsf{X}^{(1)})$ in (20) satisfies $\langle{\hat{\mathsf{K}}_{1},\mathsf{K}^{\star}}\rangle\geq k^{2}\epsilon$ with high probability. Finally, the separation assumption implies that there exist a unique $j\in[m]$ such that $\langle{\hat{\mathsf{K}}_{1},\mathsf{K}^{\star}}\rangle=\langle{\hat{\mathsf{K}}_{1},\mathsf{K}_{j}^{\star}}\rangle\geq k^{2}\epsilon$, and $\langle{\hat{\mathsf{K}}_{1},\mathsf{K}_{\ell}^{\star}}\rangle=0$, for $\ell\neq j$. Then, in the second step of the peeling algorithm, we first compute $\mathsf{X}_{ij}^{(2)}$, by setting $[\mathsf{X}_{ij}^{(2)}]_{ij}=-\infty$, for any $(i,j)\in\hat{\mathsf{K}}_{1}$, and $[\mathsf{X}_{ij}^{(2)}]_{ij}=0$, otherwise. Thus, it is clear that in the second step, $\hat{\mathsf{K}}_{2}$ cannot be attained by any set that is $k$-closed to $\hat{\mathsf{K}}_{1}$; indeed, ${\cal S}_{2}(\mathsf{K})=-\infty$, for any set $\mathsf{K}$ that is $k$-closed to $\hat{\mathsf{K}}_{1}$. Therefore, for the relevant sets in the maximization in $\hat{\mathsf{K}}_{2}$, we again have ${\cal S}_{2}(\mathsf{K})\sim{\cal N}(\lambda\langle{\mathsf{K},\mathsf{K}^{\star}}\rangle,k^{2})$. Accordingly, repeating the exact same arguments above, we obtain that $\langle{\hat{\mathsf{K}}_{2},\mathsf{K}_{j}^{\star}}\rangle\geq k^{2}\epsilon$ with high probability, for some $j\in[m]$. In the same way, we get that $\langle{\hat{\mathsf{K}}_{\ell},\mathsf{K}_{\ell}^{\star}}\rangle\geq k^{2}\epsilon$ with high probability, for any $\ell\in[m]$. The union bound, then implies that
$\mathbb{P}\left[\bigcup_{\ell=1}^{m}\left\{\langle{\hat{\mathsf{K}}_{\ell},\mathsf{K}_{\ell}^{\star}}\rangle<k^{2}\epsilon\right\}\right]\leq m\cdot\mathbb{P}\left[\langle{\hat{\mathsf{K}}_{\ell},\mathsf{K}_{\ell}^{\star}}\rangle<k^{2}\epsilon\right]\leq me^{-\Omega(n)}=o(1).$ (144)
@@ 4.6.1 Exact recovery
    We let $\mathsf{K}_{(j-1)\alpha+i}$, for $j=1,2,\ldots,m$ and $i=1,2,\ldots,\alpha$, to be defined the same as $\mathsf{K}_{0}$ but with $\mathsf{S}^{j-1}$ shifted $i$ columns to the right.
@@
$=\frac{1}{(\mathsf{M}+1)}\sum_{j=0}^{\mathsf{M}}d_{\mathsf{KL}}(\bar{\mathbb{P}}_{0}||\bar{\mathbb{P}}_{j})$ (151)
$=\frac{2km}{(\mathsf{M}+1)}\left[d_{\mathsf{KL}}\left(\mathcal{N}(\lambda,1)||\mathcal{N}(0,1)\right)+d_{\mathsf{KL}}\left(\mathcal{N}(0,1)||\mathcal{N}(\lambda,1)\right)\right]\sum_{j=0}^{\alpha}j$ (152)
$=\frac{2km}{(\mathsf{M}+1)}\frac{\alpha(1+\alpha)}{2}\left[d_{\mathsf{KL}}\left(\mathcal{N}(\lambda,1)||\mathcal{N}(0,1)\right)+d_{\mathsf{KL}}\left(\mathcal{N}(0,1)||\mathcal{N}(\lambda,1)\right)\right]$ (153)
$\leq 2k\frac{(1+\alpha)}{2}\left[d_{\mathsf{KL}}\left(\mathcal{N}(\lambda,1)||\mathcal{N}(0,1)\right)+d_{\mathsf{KL}}\left(\mathcal{N}(0,1)||\mathcal{N}(\lambda,1)\right)\right]$ (154)
$=(1+\alpha)k\lambda^{2}.$ (155)
@@
$\lambda^{2}<\frac{\frac{1}{2}\log(1+\mathsf{M})-1}{(1+\alpha)k}=\frac{\frac{1}{2}\log{(1+\alpha m)}-1}{(1+\alpha)k}.$ (156)
Finally, it is clear that there exists $\alpha_{0}\in\mathbb{N}$, such that for any $\alpha>\alpha_{0}$, the minimax error probability is at least half, if $\lambda^{2}<C/k$, for some constant $C>0$, which concludes the proof.